OpenStudy (anonymous):

Use implict differation to find the following: 6x^2-y^2=xy, (-1,3) (a) The equation of the tangent line at the indicated point on the graph. y=?

OpenStudy (helder_edwin):

this is the same as the other problems you wrote! what is it you don't understand?

OpenStudy (anonymous):

I know i tried doing it and i couldnt get it right y-3=-1(x+1)?

OpenStudy (anonymous):

or y-3=-3(x+1)?

OpenStudy (anonymous):